OpenStudy (anonymous):

Can someone help me? If a sphere has a diameter of 26 units then what is the surface area of the sphere? A.) 507 π square units B.)1,352 π square units C.)2,929 π square units D.) 676 π sqaure units

OpenStudy (anonymous):

\[A = 4 \pi r^2 = 4 \pi 13^2 = 676 \pi\]

OpenStudy (anonymous):

You just posted this.. Don't repost.

OpenStudy (anonymous):

Did you find the surface area equation?

OpenStudy (anonymous):

@ muh4ha thanks so much :)

OpenStudy (anonymous):

no problem katie - but you rly should try and understand it, so you can do it on your own. Did you get, why I plugged in 13 ?

OpenStudy (anonymous):

ni can you explain?

OpenStudy (anonymous):

Because the 'r' in the Formula is the 'radius' of the sphere - that's the distance from the center to the surface. You are given the diameter - that's twice as much (same as when looking at a circle).. So another way of writing the formula would be: \[A = 4 \pi ({d \over 2})^2\] okay? Did you get it? Any other questions?

OpenStudy (anonymous):

oh wow thanks so much :)

OpenStudy (anonymous):

you're welcome :) best thing to do, would be to try another example completely on your own and see if you can get it ;)